OpenStudy (anonymous):

Please help ASAP! verify the identity: (SinX/Sinx+CosX)=(TanX/1+TanX)

OpenStudy (anonymous):

I'd start by rewriting the right hand side. Use the following information:\[\huge tanx=\frac{sinx}{cosx}\]

OpenStudy (anonymous):

I did that. Got me to a point where it can't equal to SinX/SinX+CosX

OpenStudy (anonymous):

\[\huge \frac{ sinx }{ sinx+cosx } = \frac{ \frac{ sinx }{ cosx } }{ 1+\frac{ sinx }{ cosx } }\] \[\huge ...= \frac{ \frac{ sinx }{ cosx } }{ \frac{ cosx }{ cosx }+\frac{ sinx }{ cosx } }\]

OpenStudy (anonymous):

Since you did say hey sexy people I decided to join the party and you're really making this too hard ... \[\frac{ \tan x }{ 1+\tan x }=\frac{ \frac{ sinx}{ \cos x } }{ 1+\frac{\sin x }{ \cos x } }\] finish it by multiplying numerator and denominator by \cos{x} ...

OpenStudy (anonymous):

OOOOOOH MY GOOOD. I feel so stupid...

OpenStudy (anonymous):

Thank you guys so much, that should work out because I didn't multiple numerator and denominator by CosX

OpenStudy (anonymous):

lol Anytime! :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I have a few more questions, you guys mind helping me out again?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

..... that question is done I think @Loser66

OpenStudy (anonymous):

It's always nice having alternative solutions

OpenStudy (anonymous):

If I do it the way Loser66 did it would it count as "verification"?

OpenStudy (anonymous):

true that is of course another way

OpenStudy (anonymous):

I was just simply stating that the question had already technically been answered he didn't have to delete it..

OpenStudy (anonymous):

cyberbully

OpenStudy (anonymous):

HAHAHA aww snap

OpenStudy (anonymous):

ah now come on bib I was not calling him bad names nor did I say he was incorrect so shut your little kitty piehole

OpenStudy (anonymous):

shutting my piehole in 3..2...

OpenStudy (anonymous):

you forgot the 1....

OpenStudy (anonymous):

I started the silence at 1. Anyhow post a new question and tag us mr/ms precal

OpenStudy (anonymous):

I think mr/ms has a username bib

OpenStudy (anonymous):

@PrecalNarb PLS

OpenStudy (anonymous):

I shall. One second my sexy babes.

OpenStudy (anonymous):

lol ok little owl

OpenStudy (anonymous):

Uhh, I have a question, How do I put the question in like you guys answered this question? Like, enlarged and stuff.

OpenStudy (anonymous):

@shelbygt520 @wet ???? :O

OpenStudy (anonymous):

Uhhm bib help???

OpenStudy (anonymous):

I used Latex. \'[\huge blah blah \sinx\] remove the apostrophe and you get \[\huge blah blah \sinx\]

OpenStudy (anonymous):

I'm asking because the next question is full of squares and stuff

OpenStudy (anonymous):

\'[x^2\]

OpenStudy (anonymous):

format things like that \ [ stuff \]